OpenStudy (anonymous):

I will give medal and fan!!! Match each pair of equivalent ratios. 1. 3 to 5 2. 4 over 16 3. 36:40 4. 8 over 40 a. 9 to 10 b. 12 over 20 c. 20% d. 3:12

OpenStudy (anonymous):

anybody, please

OpenStudy (anonymous):

medal and fan

OpenStudy (mathstudent55):

Start with 3 to 5

OpenStudy (mathstudent55):

Look in the second column. You can eliminated d., since 3:5 is definitely not the same as 3:12

OpenStudy (cwrw238):

clue for number 1 multiply the 3 and 5 by a number less than 5

OpenStudy (anonymous):

i think its b

OpenStudy (mathstudent55):

Start with 1. 3 to 5 is the same as 3/5 and the same as 3:5 Now look in the second column. You can eliminate d., since 3:5 is definitely not the same as 3:12.

OpenStudy (anonymous):

@cwrw238

OpenStudy (mathstudent55):

For a, you need 9:10 If you multiply 3:5 by 3 you get 9:15, so a. is out.

OpenStudy (anonymous):

ok @mathstudent55

OpenStudy (mathstudent55):

b. is 12/20 Multiply 3/5 by 4 to get 12 on the left side. 3/5 becomes 12/20, so b. is it for 1.

OpenStudy (anonymous):

for number 2

OpenStudy (mathstudent55):

Now try 2. 4 over 16 = 4/16 = 4:16 4 is one fourth of 16. Is there any ratio in the second column in which the first number is also one fourth of the second number?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

the last answer on the second column

OpenStudy (mathstudent55):

Right 3/12 = 1/4 and 4/16 = 1/4, so 2 is d.

OpenStudy (anonymous):

I already have 3, its a

OpenStudy (anonymous):

3 goes with a

OpenStudy (mathstudent55):

Now 3. 36:40 You can write a ratio as a fraction, 36/40. Now try to reduce the fraction. Are 36 and 40 divisible by the same number? What number?

OpenStudy (mathstudent55):

Oh, I see, you're ahead of me. Yes, 3 a.

OpenStudy (anonymous):

so the last one is c

OpenStudy (mathstudent55):

36/40 divided by 4/4 is 9:10

OpenStudy (anonymous):

thank you very much

OpenStudy (mathstudent55):

Now the last one can only be one choice since you have already use the other 3 choices. But I'll still explain it. 8 to 40 can be written as 8/40 and reduced to 1/5 by dividing the numerator and denominator by 8. 1/5 as a decimal is 0.2, and as a percent it's 20%.

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (mathstudent55):

BTW, keep in mind that percent means a ratio out of 100, so 20% is the same as 20 to 100 or 20:100 or 20/100.

OpenStudy (anonymous):

ok bye

OpenStudy (mathstudent55):

Bye